Jane is 12 years older than Kate and Kate is Jane's age divided by 3. How old is each girl?
Answer by ewatrrr(24785)


Hi
Jane is 12 years older than Kate
Let x and (x+12)represent kate and jane's age respectively
Question states***
x = (x+12)/3
Solving for x
3x = x+12
2x = 12
x = 6, Kate's age. Jane is 18. (6 + 12)
CHECKING our Answer***
6 = 18/3